Find limit superior and limit inferior for the sequence (an)n∈N=((1/n)+(-1)^n)n∈N


1
Expert's answer
2021-01-29T05:10:57-0500

"\{a_n\}=\{\frac{1}{n}+(-1)^n\}"

"|a_n|=|\frac{1}{n}+(-1)^n|\leq2" for all integers "n\geq1". Hence "\{a_n\}" is bounded.

The first few terms are "\{0,\frac{3}{2},\frac{-2}{3},\frac{5}{4},\frac{-4}{5},\frac{7}{6},\cdots\}". The subsequence "\{a_{2n}\}=\{\frac{3}{2},\frac{5}{4},\frac{7}{6},\cdots\}" converges to 1 and the subsequence "\{a_{2n-1}\}=\{0,\frac{-2}{3},\frac{-4}{3},\cdots\}" converges to -1. Hence "S=\{-1,1\}"

"\liminf a_n=\inf(S)=-1\\
\limsup a_n=\sup(S)=1"